# Enhancing Inertial Hand based HAR through Joint Representation of Language, Pose and Synthetic IMUs

### v1

@@ Abstract.
HAR, Sensor simulation, Multi-modal contrastive learning, Pretraining
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 2024 International Symposium on Wearable Computers; October 5–9, 2023; Melbourne, Austrailiabooktitle: Proceedings of the 2024 International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, Australiasubmissionid: 8562ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Modeling and simulationccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
@@ 1. Introduction
To utilize the monocular video data and synthesized IMU data for HAR in complex domains faces two challenges: (1) Low-quality synthetic IMU data. Existing approaches for generating synthetic IMU data from videos, which require kinetic detection and motion capture, may introduce errors due to lighting, body shapes, and occlusion, thus limiting IMU generation for complex activities. Subtle motions, especially on the wrist, can be challenging to simulate from monocular video due to the wrist’s degrees of freedom and its relatively small size in pixels. For instance, as depicted in Figure 1, the synthetic IMU data for the “pick-up paper” activity exhibits considerable fluctuations and errors. While simulated data can capture part of the signal characteristics, there is clearly a simulation gap. (2) Directly training on synthetic IMU for HAR results in poor performance for fine-grained activities. Due to inherent errors mentioned in (1), synthetic IMU data shows heightened sensitivity to activities with smaller amplitudes or greater complexity. Thus models trained with synthetic IMU data may not demonstrate superior performance in HAR.
@@
(3) Through extensive experimentation to demonstrate the effectiveness of the proposed approach over other state-of-the-art IMU simulation pipelines through IMU-based HAR.
@@ 2. Related work
2.1. IMU Simulation
Various approaches have used other widely available modalities like text and videos to generate the inertial data using 3D simulation and generative models to generate IMU data and use it along with some real data to train HAR models.
@@ 2.1. IMU Simulation
Video to IMU models. In (Rey et al., 2019), virtual IMU data was generated from video data, initially extracting 2D poses via OpenPose (Cao et al., 2017), followed by the application of a regression model to translate sequences of poses into IMU data. However, training the regression model requires pre-collected pairs of pose and IMU data. Kwon et al. introduced IMUSim (Young et al., 2011) to directly obtain IMU data from sequences of poses. Nevertheless, the IMU data produced by the above methods often contain considerable errors compared to actual IMU data, thus limiting the utility of using generated data for complex HAR (Leng et al., 2023a, 2024). Santhalingam et al. (Santhalingam et al., 2023) proposed a bi-directional LSTM-based model to calibrate anomalous poses with the assistance of surrounding poses, aiming to reduce the amount of incorrectly generated IMU data. However, it does not entirely prevent errors in the translation from pose to generated IMU data. Liu et al. (Liu et al., 2021) proposed a CNN-based model to automatically calibrate errors between generated and real IMU data. However, the performance of this approach is influenced by the characteristics of the training data.
Text to IMU models. Many studies (Leng et al., 2023b, 2024) have proposed the utilization of textual descriptions converted into human poses for generating IMU data. Various pretrained pose syntheses models, such as MotionGPT (Jiang et al., 2024), T2M-GPT (Zhang et al., 2023), and MotionDiffuse (Zhang et al., 2022), have been introduced into synthetic human poses, which are then integrated with IMU generation models to obtain synthesized IMU data. However, obtaining the pose is infeasible if the type of pose has never been encountered during the training of the pose synthesis models. Thus, generating IMU data from video has become the primary approach, especially in professional domains.
2.2. Representation Learning
@@
Contrastive learning has garnered attention in recent years due to its efficacy in learning representations from data originating from diverse domains. For instance, CLIP (Radford et al., 2021) learns visual and text representations using paired images and text, achieving outstanding generalization performance for downstream tasks. Moon et al. (Moon et al., 2023) proposed a multi-modal contrastive framework and a pretraining approach to align IMU data with text and video data, thereby projecting multi-modal data representations into a joint space. Yang et al. (Yang et al., 2024) further enhanced the contrastive framework between text and IMU data by introducing a hierarchical temporal transformer to align important representations. Different from the aforementioned multi-modal contrastive approaches, our method exclusively utilizes video as input and generates both human pose and highly accurate synthesized IMU data as inputs for the contrastive model. This enables the creation of improved representation spaces for video and IMU data.
@@ 3.2. $\text{Multi}^{3}$et for Pretraining
The Pose2IMU regression block consists of a Pose encoder and a Pose2IMU decoder, which has a CNN architecture with three ConvTranspose and Unpooling layers along with Batch normalization and Dropout blocks. Since activities in real scenarios typically involve fine-grained motions, the Pose2IMU decoder is designed to guarantee that the input pose encoder encompasses the requisite features by reconstructing the IMU data from the encoder. For the Pose2IMU decoder, the input is $e_{p}$, and the output is the predicted result of the corresponding synthetic IMU data, denoted as $X_{p}$. The decoder architecture is based on PSN from PresSim (Ray et al., 2023).
@@
Similar to the Pose2IMU regression block, the IMU reconstruction block comprises two IMU encoders and an IMU reconstructor. This IMU reconstructor features an identical CNN architecture with Section 3.2 that takes the concatenation of $[e_{sl},e_{sr}]$ and uses one linear layer to map it back from (256, 2) to 256 vectors to reconstruct one frame input IMU instance. The predicted IMU data is denoted as $X_{s}$.
@@
(3) $\text{InfoNCE}(q,k)=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{e^{s(q_{i},k_{i})}}{e^{s(q_{i},k_{i})}+\sum_{j=1,j\neq i}^{N}e^{s(q_{i},k_{j})}}\right)$
where $N$ is the batch size, $q_{i}$ and $k_{i}$ are the representations of the $i$-th data sample under two different augmentations, $s(q_{i},k_{i})$ is the cosine similarity score between $q_{i}$ and $k_{i}$, normalized by the temperature, and $e$ denotes the exponential function.
The cosine similarity score $s(q_{i},k_{i})$ is computed as:
@@ 4.1. Dataset and Training Details
We utilized two types of datasets (1) Large video datasets with rich hand activity representations for pertaining, and (2) Target inertial HAR datasets with wrist IMUs. To maintain consistency, all video data were resampled to 60 frames per second.
How2Sign Dataset (Duarte et al., 2021). This dataset contains more than 80 hours of sign language videos and corresponding transcripts, which provide rich information on hand and wrist movements. This dataset is applied only for pre-training.
GRAB Dataset (Taheri et al., 2020). This dataset contains approximately 4 hours of MoCap of the entire body of subjects grabbing everyday objects. It comprises data from 10 subjects engaging with 51 different everyday objects. This dataset is applied only for pre-training.
OpenPack Dataset(Yoshimura et al., 2024). Acceleration data from both the left and right wrists of 5 workers were collected using an Empatica E4 wristband with a sampling rate of 30Hz while they performed a packaging task comprising 11 classes of activities. The workers’ activities were also recorded on video to serve as ground truth.
MM-Fit Dataset(Strömbäck et al., 2020). This dataset comprises data from 10 subjects engaging in 10 gym exercises. IMU data from Mobvoi TicWatch Pro devices, sampled at 100Hz, capture detailed movement information from the participants’ wrists. Additionally, RGB data captured at 30Hz offers visual context for the performed exercises.
The model is written in Pytorch and trained on an Ubuntu system with Nvidia A6000 ada GPU and AMD Ryzen 7900X CPU. We used Adam optimizer for training. Both pretraining and downstream training are done for 300 epochs with the early stopping of patience 20 and a learning rate scheduler of patience 10 starting with 0.001.
@@ 4.2. Quantitative Results
Table 1. Weighted Macro F1-score for OpenPack dataset
@@
DCL (Real) 45.80 $\pm$ 1.97 49.48 $\pm$ 2.23
DCL (Real + Synthetic Vi2IMU) 35.41 $\pm$ 1.12 37.86 $\pm$ 1.06
DCL (Real + Synthetic IMUTube) 36.98 $\pm$ 2.23 48.66 $\pm$ 0.63
Base (Real) 31.43 $\pm$ 0.25 53.84 $\pm$ 0.15
Base (Real + Synthetic Vi2IMU) 40.03 $\pm$ 0.47 32.60 $\pm$ 0.54
Base (Real + Synthetic IMUTube) 24.96 $\pm$ 0.87 31.88 $\pm$ 1.69
IMU Reconstruction (how2sign:frozen) 19.99 $\pm$ 0.32 30.53 $\pm$ 0.27
IMU Reconstruction (how2sign:not frozen) 31.79 $\pm$ 0.41 53.44 $\pm$ 0.36
Contrastive pretrain (how2sign:frozen) 20.64 $\pm$ 0.22 30.18 $\pm$ 0.29
Contrastive pretrain (how2sign:not frozen) 39.86 $\pm$ 0.34 59.11 $\pm$ 0.40
IMU Reconstruction (GRAB:frozen) 20.48 $\pm$ 0.18 30.22 $\pm$ 0.26
IMU Reconstruction (GRAB:not frozen) 31.41 $\pm$ 0.19 54.00 $\pm$ 0.30
Contrastive pretrain (GRAB:frozen) 21.19 $\pm$ 0.20 32.82 $\pm$ 0.33
Contrastive pretrain (GRAB:not frozen) 33.10 $\pm$ 0.28 54.91 $\pm$ 0.37
$\text{Multi}^{3}$et (how2sign:frozen) 20.53 $\pm$ 0.23 30.88 $\pm$ 0.31
$\text{Multi}^{3}$et (how2sign:not frozen) 42.32 $\pm$ 0.13 61.74 $\pm$ 0.22
$\text{Multi}^{3}$et (GRAB:frozen) 20.77 $\pm$ 0.21 32.96 $\pm$ 0.35
$\text{Multi}^{3}$et (GRAB:not frozen) 34.08 $\pm$ 0.38 55.12 $\pm$ 0.41
$\text{Multi}^{3}$et (Both: frozen) 25.13 $\pm$ 0.24 38.29 $\pm$ 0.39
$\text{Multi}^{3}$et (Both:not frozen) 39.38 $\pm$ 0.37 60.23 $\pm$ 0.35
Table 2. Weighted Macro F1-score for MM-Fit dataset
Model Left wrist Both wrists
DCL (Real) 90.66 $\pm$ 0.41 90.59 $\pm$ 0.53
DCL (Real + Synthetic Vi2IMU) 90.39 $\pm$ 1.06 89.73$\pm$ 0.49
DCL (Real + Synthetic IMUTube) 89.73 $\pm$ 0.81 89.95 $\pm$ 0.68
Base (Real) 80.86 $\pm$ 0.35 90.11 $\pm$ 0.55
Base (Real + Synthetic Vi2IMU) 89.00 $\pm$ 0.10 86.39 $\pm$ 0.15
Base (Real + Synthetic IMUTube) 87.24 $\pm$ 0.17 89.19 $\pm$ 0.92
IMU reconstruction (how2sign:frozen) 28.76 $\pm$ 0.62 35.48 $\pm$ 0.41
IMU reconstruction (how2sign:not frozen) 80.91 $\pm$ 0.52 90.08 $\pm$ 0.33
Contrastive pretrain (how2sign:frozen) 35.78 $\pm$ 0.47 40.01 $\pm$ 0.37
Contrastive pretrain (how2sign:not frozen) 87.23 $\pm$ 0.59 91.85 $\pm$ 0.69
IMU reconstruction (GRAB:frozen) 28.89 $\pm$ 0.29 35.66 $\pm$ 0.51
IMU reconstruction (GRAB:not frozen) 80.89 $\pm$ 0.48 90.17 $\pm$ 0.66
Contrastive pretrain (GRAB:frozen) 37.18 $\pm$ 0.40 44.32 $\pm$ 0.60
Contrastive pretrain (GRAB:not frozen) 85.03 $\pm$ 0.57 90.17 $\pm$ 0.42
$\text{Multi}^{3}$et (how2sign:frozen) 37.03 $\pm$ 0.31 41.86 $\pm$ 0.55
$\text{Multi}^{3}$et (how2sign:not frozen) 89.49 $\pm$ 0.67 93.02 $\pm$ 0.71
$\text{Multi}^{3}$et (GRAB:frozen) 38.41 $\pm$ 0.38 45.93 $\pm$ 0.53
$\text{Multi}^{3}$et (GRAB:not frozen) 87.71 $\pm$ 0.46 90.98 $\pm$ 0.61
$\text{Multi}^{3}$et (Both:frozen) 40.33 $\pm$ 0.49 46.57 $\pm$ 0.70
$\text{Multi}^{3}$et (Both:not frozen) 88.18 $\pm$ 0.63 92.19 $\pm$ 0.68
@@
In this section, we compare the proposed approach with synthetic IMU generated by different approaches and directly used for HAR. 5 times leave-one-user-out experiments were conducted with different random seeds.
DCL (Real): Only real IMU data of the target dataset is used for training, the HAR model is DeepConvLSTM (DCL) (Ordóñez and Roggen, 2016), which is widely used for IMU-based HAR tasks.
DCL (Real+Synthetic Vi2IMU): Both synthetic and real IMU data from the target dataset are used for training. The synthetic IMU is generated by Vi2IMU (Santhalingam et al., 2023), which takes 2D and 3D pose and orientation series as input, and an LSTM-based model is used to predict the virtual IMU. Note that as the orientation is not available in the paper, we utilize OSX (Lin et al., 2023) to output the approximate orientations.
DCL (Real+Synthetic IMUTube): Both synthetic and real IMU data from the target dataset are used for training. The synthetic IMU is created using IMUTube (Kwon et al., 2020), starting with extracting 2D skeletal poses from videos using AlphaPose (Fang et al., 2022), then mapping the 2D poses to 3D using VideoPose3D (Pavllo et al., 2019), and using IMUSim to generate synthetic IMU data for specific body joints. Finally the simulated IMU is calibrated using some real IMU data from the target set to have similar range of variability.
Base (Real), (Real+Synthetic Vi2IMU), and (Real+Synthetic IMUTube): We do the downstream training without any pretraining. The IMU encoder weights are initialized randomly. We use real IMU data, real and synthetic IMU data generated by Vi2IMU and IMUTube, to train the model.
IMU reconstruction (frozen) and (not frozen): Only the IMU reconstruction model is applied for pretraining. ”Frozen” indicates that the learned weights of the IMU encoder are frozen, and only the classifier parameters are optimized. ”Not frozen” means that the IMU encoder is frozen for N epochs until the loss stops decreasing and reaches the patience P for the first time, after which it is unfrozen while the classifier is learning.
Contrastive pretrain: Only the multi-modal contrastive model is applied for pretraining.
Multi3Net: The proposed approach utilizes multi-task pretraining to create a better joint representation for downstream classification.
Tables 1 and 2 display the HAR performance for the OpenPack and MM-Fit datasets, respectively. Overall, the proposed approach outperforms the methods that utilize real IMU or real and synthetic IMU data directly for HAR on both wrists. Specifically, the proposed method demonstrates an 7.90 % improvement for OpenPack with both wrists and 2.43 % improvement for MM-Fit with both wrists. This indicates that useful information can be extracted from extensive datasets with activities not present in the target dataset, effectively addressing the lack of data issue. However, the performance of the proposed approach with the left wrist is poorer than DCL (Real), this may be due to the fact that information obtained from a single body part using the proposed method may not be as accurate, whereas obtaining data from multiple body parts can enhance the HAR performance. Besides, the macro F1-score exhibited poorer performance in many scenarios when utilizing synthetic IMU generated by Vi2IMU. This discrepancy may be attributed to errors in calculating the joint orientation, indicating that the quality of synthetic IMU data significantly impacts HAR performance for complex activities.

### v2

@@ Abstract.
HAR, Sensor simulation, Multi-modal learning, Pretraining
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM International Symposium on Wearable Computers; October 5–9, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2024 ACM International Symposium on Wearable Computers (ISWC ’24), October 5–9, 2024, Melbourne, VIC, Australiadoi: 10.1145/3675095.3676609isbn: 979-8-4007-1059-9/24/10submissionid: 8562ccs: Computing methodologies Learning latent representationsccs: Computing methodologies Modeling and simulationccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methods
@@ 1. Introduction
To utilize the monocular video data and synthesized IMU data for HAR in complex domains one faces two challenges: (1) Low-quality synthetic IMU data. Existing approaches for generating synthetic IMU data from videos, which require kinetic detection and motion capture, may introduce errors due to lighting, body shapes, and occlusion, thus limiting IMU generation for complex activities. Subtle motions, especially on the wrist, can be challenging to simulate from monocular video due to the wrist’s degrees of freedom and its relatively small size in pixels. For instance, as depicted in Figure 1, the synthetic IMU data for the “pick-up paper” activity exhibits considerable fluctuations and errors. While simulated data can capture part of the signal characteristics, there is clearly a simulation gap. (2) Directly training on synthetic IMU for HAR results in poor performance for fine-grained activities. Due to inherent errors mentioned in (1), synthetic IMU data shows heightened sensitivity to activities with smaller amplitudes or greater complexity. Thus models trained with synthetic IMU data may not demonstrate superior performance in HAR.
@@
(3) Through extensive experimentation we demonstrate the effectiveness of the proposed approach over other state-of-the-art IMU simulation pipelines through IMU-based HAR.
@@ 2. Related work
IMU Simulation: Various approaches have used other widely available modalities like text and videos to generate the inertial data using 3D simulation and generative models to generate IMU data and use it along with some real data to train HAR models. In Video to IMU models, virtual IMU data was generated from video data (Rey et al., 2019), initially extracting 2D poses via OpenPose (Cao et al., 2017), followed by the application of a regression model to translate sequences of poses into IMU data. However, training the regression model requires pre-collected pairs of pose and IMU data. Young et al. introduced IMUSim (Young et al., 2011) to directly obtain IMU data from sequences of poses. Nevertheless, the IMU data produced by the above methods often contain considerable errors compared to actual IMU data, thus limiting the utility of using generated data for complex HAR (Leng et al., 2023a, 2024). Santhalingam et al. (Santhalingam et al., 2023) proposed a bi-directional LSTM-based model to calibrate anomalous poses with the assistance of surrounding poses, aiming to reduce the amount of incorrectly generated IMU data. However, it does not entirely prevent errors in the translation from pose to generated IMU data. Liu et al. (Liu et al., 2021) proposed a CNN-based model to automatically calibrate errors between generated and real IMU data. However, the performance of this approach is influenced by the characteristics of the training data. Similarly many Text to IMU models(Leng et al., 2023b, 2024) have proposed the utilization of textual descriptions converted into human poses for generating IMU data. Various pretrained pose syntheses models, such as MotionGPT (Jiang et al., 2024), T2M-GPT (Zhang et al., 2023), and MotionDiffuse (Zhang et al., 2022), have been introduced into synthetic human poses, which are then integrated with IMU generation models to obtain synthesized IMU data. However, obtaining the pose is infeasible if the type of pose has never been encountered during the training of the pose synthesis models. Thus, generating IMU data from video has become the primary approach, especially in professional domains.
@@
Representation Learning: Contrastive learning has garnered attention in recent years due to its efficacy in learning representations from data originating from diverse domains. For instance, CLIP (Radford et al., 2021) learns visual and text representations using paired images and text, achieving outstanding generalization performance for downstream tasks. Moon et al. (Moon et al., 2023) proposed a multi-modal contrastive framework and a pretraining approach to align IMU data with text and video data, thereby projecting multi-modal data representations into a joint space. Yang et al. (Yang et al., 2024) further enhanced the contrastive framework between text and IMU data by introducing a hierarchical temporal transformer to align important representations. Different from the aforementioned multi-modal contrastive approaches, our method exclusively utilizes video as input and generates both human pose and highly accurate synthesized IMU data as inputs for the contrastive model. This enables the creation of improved representation spaces for video and IMU data.
@@ 3.2. $\text{Multi}^{3}$et for Pretraining
The Pose2IMU regression block consists of a Pose encoder and a Pose2IMU decoder, which has a CNN architecture with three ConvTranspose and Unpooling layers along with Batch normalization and Dropout blocks. Since activities in real scenarios typically involve fine-grained motions, the Pose2IMU decoder is designed to guarantee that the input pose encoder encompasses the required features by reconstructing the IMU data from the encoder. For the Pose2IMU decoder, the input is $e_{p}$, and the output is the predicted result of the corresponding synthetic IMU data, denoted as $X_{p}$. The decoder architecture is based on PSN from PresSim (Ray et al., 2023).
@@
Similar to the Pose2IMU regression block, the IMU reconstruction block comprises two IMU encoders and an IMU reconstructor. This IMU reconstructor features an identical CNN architecture to that of Pose2IMU but that takes the concatenation of $[e_{sl},e_{sr}]$ and uses one linear layer to map it back from (256, 2) to 256 vectors to reconstruct one frame input IMU instance. The predicted IMU data is denoted as $X_{s}$.
@@
(3) $\text{InfoNCE}(q,k)=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{e^{s(q_{i},k_{i})/\tau}}{e^{s(q_{i},k_{i})/\tau}+\sum_{j=1,j\neq i}^{N}e^{s(q_{i},k_{j})/\tau}}\right)$
where $N$ is the batch size, $q_{i}$ and $k_{i}$ are the representations of the $i$-th data sample under two different augmentations, $s(q_{i},k_{i})$ is the cosine similarity score between $q_{i}$ and $k_{i}$, normalized by the temperature $\tau$ of 0.07. The cosine similarity score $s(q_{i},k_{i})$ is computed as:
@@ 4.1. Dataset and Training Details
We utilized two types of datasets (1) Large video datasets with rich hand activity representations for pretraining, and (2) Target inertial HAR datasets with wrist IMUs. To maintain consistency, all video data were resampled to 60 frames per second.
How2Sign Dataset (Duarte et al., 2021) contains more than 80 hours of sign language videos and corresponding transcripts, which provide rich information on hand and wrist movements. This dataset is applied only for pretraining. GRAB Dataset (Taheri et al., 2020) contains approximately 4 hours of MoCap of the entire body of subjects grabbing everyday objects. It comprises data from 10 subjects engaging with 51 different everyday objects. This dataset is applied only for pretraining. OpenPack Dataset(Yoshimura et al., 2024) contains acceleration data from both the left and right wrists of 5 workers were collected using an Empatica E4 wristband with a sampling rate of 30Hz while they performed a packaging task comprising 11 classes of activities. The workers’ activities were also recorded on video to serve as ground truth. MM-Fit Dataset(Strömbäck et al., 2020) comprises data from 10 subjects engaging in 10 gym exercises. IMU data from Mobvoi TicWatch Pro devices, sampled at 100Hz, capture detailed movement information from the participants’ wrists. Additionally, RGB data captured at 30Hz offers visual context for the performed exercises.
@@ 4.2. Quantitative Results
Table 1. Macro F1-score for different datasets.
@@
OpenPack Dataset
DCL (Real) 43.25 $\pm$ 0.81 43.09 $\pm$ 0.50
DCL (Real + Synthetic Vi2IMU) 32.71 $\pm$ 0.75 33.50 $\pm$ 0.68
DCL (Real + Synthetic IMUTube) 42.48 $\pm$ 1.56 41.34 $\pm$ 1.48
Base (Real) 33.79 $\pm$ 0.39 42.26 $\pm$ 0.25
Base (Real + Synthetic Vi2IMU) 39.19 $\pm$ 0.18 38.28 $\pm$ 0.39
Base (Real + Synthetic IMUTube) 35.28 $\pm$ 0.74 40.29 $\pm$ 0.89
IMU Reconstruction (how2sign:frozen) 33.21 $\pm$ 0.53 41.19 $\pm$ 0.33
IMU Reconstruction (how2sign:not frozen) 39.71 $\pm$ 0.24 48.71 $\pm$ 0.45
Contrastive pretrain (how2sign:frozen) 39.39 $\pm$ 0.37 53.65 $\pm$ 0.18
Contrastive pretrain (how2sign:not frozen) 45.27 $\pm$ 0.18 58.24 $\pm$ 0.26
IMU Reconstruction (GRAB:frozen) 31.36 $\pm$ 0.13 39.14 $\pm$ 0.19
IMU Reconstruction (GRAB:not frozen) 37.33 $\pm$ 0.41 46.18 $\pm$ 0.20
Contrastive pretrain (GRAB:frozen) 40.21 $\pm$ 0.23 53.78 $\pm$ 0.53
Contrastive pretrain (GRAB:not frozen) 44.11 $\pm$ 0.26 57.17 $\pm$ 0.38
$\text{Multi}^{3}$et (how2sign:frozen) 40.41 $\pm$ 0.17 54.17 $\pm$ 0.28
$\text{Multi}^{3}$et (how2sign:not frozen) 47.32 $\pm$ 0.13 59.83 $\pm$ 0.27
$\text{Multi}^{3}$et (GRAB:frozen) 41.22 $\pm$ 0.16 55.07 $\pm$ 0.34
$\text{Multi}^{3}$et (GRAB:not frozen) 45.18 $\pm$ 0.28 58.28 $\pm$ 0.28
$\text{Multi}^{3}$et (Both: frozen) 41.88 $\pm$ 0.28 56.36 $\pm$ 0.16
$\text{Multi}^{3}$et (Both:not frozen) 48.39 $\pm$ 0.18 61.07 $\pm$ 0.39
MM-Fit Dataset
DCL (Real) 75.48 $\pm$ 2.53 75.79 $\pm$ 2.02
DCL (Real + Synthetic Vi2IMU) 74.61 $\pm$ 1.78 73.73 $\pm$ 2.62
DCL (Real + Synthetic IMUTube) 75.63 $\pm$ 1.56 75.97 $\pm$ 2.35
Base (Real) 85.18 $\pm$ 0.31 88.13 $\pm$ 0.57
Base (Real + Synthetic Vi2IMU) 87.85 $\pm$ 0.48 86.59 $\pm$ 0.15
Base (Real + Synthetic IMUTube) 83.37 $\pm$ 0.26 88.86 $\pm$ 0.25
IMU reconstruction (how2sign:frozen) 75.63 $\pm$ 0.18 78.37 $\pm$ 0.41
IMU reconstruction (how2sign:not frozen) 82.74 $\pm$ 0.38 86.56 $\pm$ 0.22
Contrastive pretrain (how2sign:frozen) 80.66 $\pm$ 0.61 84.52 $\pm$ 0.33
Contrastive pretrain (how2sign:not frozen) 89.19 $\pm$ 0.74 93.45 $\pm$ 0.71
IMU reconstruction (GRAB:frozen) 77.19 $\pm$ 0.34 82.11 $\pm$ 0.68
IMU reconstruction (GRAB:not frozen) 83.53 $\pm$ 0.64 87.15 $\pm$ 0.38
Contrastive pretrain (GRAB:frozen) 80.47 $\pm$ 0.49 86.64 $\pm$ 0.53
Contrastive pretrain (GRAB:not frozen) 88.28 $\pm$ 0.61 90.37 $\pm$ 0.18
$\text{Multi}^{3}$et (how2sign:frozen) 80.55 $\pm$ 0.18 86.38 $\pm$ 0.82
$\text{Multi}^{3}$et (how2sign:not frozen) 91.03 $\pm$ 0.13 93.81 $\pm$ 0.29
$\text{Multi}^{3}$et (GRAB:frozen) 82.33 $\pm$ 0.17 87.59 $\pm$ 0.21
$\text{Multi}^{3}$et (GRAB:not frozen) 89.72 $\pm$ 0.17 91.98 $\pm$ 0.20
$\text{Multi}^{3}$et (Both:frozen) 81.37 $\pm$ 0.81 86.28 $\pm$ 0.26
$\text{Multi}^{3}$et (Both:not frozen) 91.15 $\pm$ 0.26 93.44 $\pm$ 0.07
@@
In this Section, we compare the proposed approach with synthetic IMU generated by different approaches and directly used for HAR. We conducted leave-one-user-out experiments that were repeated 5 times with different random seeds.
For DCL (Real) only real IMU data of the target dataset is used for training, the HAR model is DeepConvLSTM (DCL) (Ordóñez and Roggen, 2016), which is widely used for IMU-based HAR tasks. For DCL (Real+Synthetic Vi2IMU) both synthetic and real IMU data from the target dataset are used for training. The synthetic IMU is generated by Vi2IMU (Santhalingam et al., 2023), which takes 2D and 3D pose and orientation series as input, and an LSTM-based model is used to predict the virtual IMU. Note that as the orientation is not available in the paper, we utilize OSX (Lin et al., 2023) to output the approximate orientations. For DCL (Real+Synthetic IMUTube) both synthetic and real IMU data from the target dataset are used for training. The synthetic IMU is created using IMUTube (Kwon et al., 2020), starting with extracting 2D skeletal poses from videos using AlphaPose (Fang et al., 2022), then mapping the 2D poses to 3D using VideoPose3D (Pavllo et al., 2019), and using IMUSim to generate synthetic IMU data for specific body joints. Finally the simulated IMU is calibrated using some real IMU data from the target set to have similar range of variability. For Base (Real), (Real+Synthetic Vi2IMU), and (Real+Synthetic IMUTube) we do the downstream training without any pretraining. The IMU encoder weights are initialized randomly. We use real IMU data, real and synthetic IMU data generated by Vi2IMU and IMUTube, to train the model. In case of IMU reconstruction (frozen) and (not frozen) only the IMU reconstruction model is applied for pretraining where ”frozen” indicates that the learned weights of the IMU encoder are frozen, while ”not frozen” means that the IMU encoder is frozen for N epochs until the loss stops decreasing and reaches the patience P for the first time, after which it is unfrozen while the classifier is learning. In Contrastive pretrain only the multi-modal contrastive model is applied for pretraining. Multi3Net is the proposed approach that utilizes multi-task pretraining to create a better joint representation for downstream classification. Tables 1 display the HAR performance for the OpenPack and MM-Fit datasets, respectively. Overall the proposed approach outperforms the methods that directly utilize real IMU or real and synthetic IMU data for HAR with one or both wrists. Specifically, the proposed method demonstrates an 18.81% macro F1 improvement for OpenPack with one wrist and 14.6% for both wrists. The gain in Macro F1 for MM-Fit with one wrist is of 5.96% while with two wrists it is of 5.31%. This improvement is significant when compared to synthetic IMU generated approaches, which in some cases decrease the performance. This discrepancy may be attributed to errors in calculating the joint orientation, indicating that the quality of synthetic IMU data significantly impacts HAR performance for complex activities.